# Distributional and mean Li–Yorke chaos for weighted shifts on Fréchet sequence spaces

### 02 Mar 2026 v1

@@ 4 Mean Li–Yorke chaos
$\liminf_{n\to\infty}\frac{1}{n}\sum_{k=0}^{n-1}d\big(f^{k}(x),f^{k}(y)\big)=0\quad\text{and}\quad\limsup_{n\to\infty}\frac{1}{n}\sum_{k=0}^{n-1}d\big(f^{k}(x),f^{k}(y)\big)>0.$
The function $f$ is said to be mean Li–Yorke chaotic if there exists an uncountable set $S\subset X$ such that every distinct pair $(x,y)\in S\times S$ is a mean Li–Yorke pair.
From now on, given a Fréchet space $X$ endowed with a family of seminorms $\bigl(\|\cdot\|_{n}\bigr)_{n\in\mathbb{N}}$, we will always consider the compatible metric $d$ to be the one given in (1). To prove our main result in this section, we need the following definitions and lemmas from [14].
@@
Let $X$ be a Fréchet space endowed with an increasing sequence $(\|\cdot\|_{k})_{k\in\mathbb{N}}$ of seminorms and with the compatible metric $d$. Let $T:X\to X$ be a continuous linear operator and let $m\in\mathbb{N}$.
@@
    A vector $x\in X$ is called absolutely mean semi-irregular if
@@
    A vector $x\in X$ is called absolutely mean $m$-irregular if
Definition 25.
Let $T:X\to X$ be a continuous linear operator on a Fréchet space $X$ with the compatible metric $d$. The mean proximal cell of 0 is the set
$\text{MProx}(T,0):=\left\{x\in X:\liminf_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d(T^{k}x,0)=0\right\}.$
@@
[14, Proposition 4.25] Let $T:X\to X$ be a continuous linear operator on a Fréchet space $X$, endowed with an increasing sequence $(\left\|\cdot\right\|_{k})_{k\in\mathbb{N}}$ of seminorms and with the compatible metric $d$. Fix $m\in\mathbb{N}$. Then, the following statements are equivalent:
  • a)
    There exists $x\in X$, such that
    $\limsup_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}\left\|T^{k}x\right\|_{m}=\infty.$
  • b)
    There exist a sequence $(y_{k})_{k\in\mathbb{N}}\subset X$ and an increasing sequence $(N_{k})_{k\in\mathbb{N}}$ of positive integers, such that $y_{k}\to 0$, when $k\to\infty$, and
    $\inf_{k\in\mathbb{N}}\frac{1}{N_{k}}\sum_{j=1}^{N_{k}}\left\|T^{j}y_{k}\right\|_{m}>0.$
@@
[14, Theorem 4.27] Let $T:X\to X$ be a continuous linear operator on a Fréchet space $X$ with the compatible metric $d$. Then, we have the implications $(1)\Longleftrightarrow(2)\Longrightarrow(3)$ where:
  1. (1)
@@
  3. (3)
    $T$ admits a residual set of absolutely mean $m$-irregular vectors for some $m\in\mathbb{N}$.
Lemma 28.
  • a)
    $T$ is mean Li–Yorke chaotic.
  • b)
    $T$ admits an absolutely mean semi-irregular vector.
Lemma 29.
@@
The following lemma is a straightforward exercise, and we leave its proof to the reader.
@@
Let $X$ be a Fréchet sequence space over $\mathbb{Z}$, endowed with an increasing sequence $(\left\|\cdot\right\|_{n})_{n\in\mathbb{N}}$ of seminorms, with the compatible metric $d$. Suppose that the sequence $(e_{n})_{n\in\mathbb{Z}}$ of canonical vectors is a basis. Suppose that the bilateral weighted backward shift $B_{w}$, with nonzero weights $w:=(w_{n})_{n\in\mathbb{Z}}$, is well-defined and continuous on $X$. Then the following statements are equivalent:
@@
    $\liminf_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d(w_{j-k}\cdots w_{j-1}e_{j-k},0)=0,$ for some $j\in\mathbb{Z}$.
@@
Let $X$ be a Fréchet sequence space over $\mathbb{Z}$, endowed with an increasing sequence $(\left\|\cdot\right\|_{n})_{n\in\mathbb{N}}$ of seminorms and with the compatible metric $d$. Suppose that the sequence $(e_{n})_{n\in\mathbb{Z}}$ of canonical vectors is a basis and that condition (C) holds. Suppose that the bilateral weighted backward shift $B_{w}$, with nonzero weights $w:=(w_{n})_{n\in\mathbb{Z}}$, is well-defined and continuous on $X$. Then $B_{w}$ is mean Li–Yorke chaotic if and only if the following conditions hold:
@@
    $\liminf_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d(w_{-k}\cdots w_{-1}e_{-k},0)=0.$
@@
($\Rightarrow$): Suppose that $B_{w}$ is mean Li-Yorke chaotic. Therefore, by Lemma 28, $B_{w}$ admits an absolutely mean semi-irregular vector $y:=(y_{j})_{j\in\mathbb{Z}}$, then
$\liminf_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d((B_{w})^{k}(y),0)=0.$ (18)
Therefore, by Lemma 30, condition (A) holds. Since $y$ is an absolutely mean semi-irregular vector, Lemma 29 ensures the existence of $m\in\mathbb{N}$ and $x=(x_{j})_{j\in\mathbb{Z}}\in X$ such that
@@
Thus, by (20) and (21), we conclude condition (B).
($\Leftarrow$): By condition (A) and Lemma 30, we have that
If there exists $j_{0}\in\mathbb{Z}$ such that $\limsup_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d((B_{w})^{k}(e_{j_{0}}),0)>0$, then $B_{w}$ is mean Li-Yorke chaotic, given that $e_{j_{0}}$ is an absolutely mean semi-irregular vector. Otherwise, we have that
Thus, using the facts: $d(x_{1}+x_{2},0)\leq d(x_{1},0)+d(x_{2},0)$ and $d(\lambda x_{1},0)\leq(1+|\lambda|)d(x_{1},0)$ for all $\lambda\in\mathbb{K}$ and for all $x_{1},x_{2}\in X$, we obtain that $c_{00}(\mathbb{Z})\subset\text{MProx}(B_{w},0)$. Therefore, $\text{MProx}(B_{w},0)$ is dense. By condition (B) and Lemma 26 (applied to $y_{k}:=\frac{\sum_{j=-r}^{r}b_{j,k}e_{j}}{k\left\|\sum_{j=-r}^{r}b_{j,k}e_{j}\right\|_{p(k)}}$, $k\in\mathbb{N}$), there exists a vector $x\in X$ such that
$\limsup_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}\left\|(B_{w})^{k}(x)\right\|_{m}=\infty\quad\text{and, consequently,}\quad\limsup_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d((B_{w})^{k}(x),0)>0.$ (22)
Using (22) and the fact that $\text{MProx}(B_{w},0)$ is dense, by Lemma 27 we conclude that $B_{w}$ admits a dense set of absolutely mean semi-irregular vectors. Therefore, by Lemma 28, $B_{w}$ is mean Li-Yorke chaotic. ∎
Remark 32.
$\lim_{n\to\infty}(B_{w})^{n}(e_{i})=0,\quad\text{for all }i\in\mathbb{N},$
@@
Let $X$ be a Fréchet sequence space, endowed with an increasing sequence $(\left\|\cdot\right\|_{n})_{n\in\mathbb{N}}$ of seminorms and with the compatible metric $d$. Suppose that the sequence $(e_{n})_{n\in\mathbb{N}}$ of canonical vectors is a basis. Suppose that the unilateral weighted backward shift $B_{w}$, with nonzero weights $w:=(w_{n})_{n\in\mathbb{N}}$, is well-defined and continuous on $X$. Then $B_{w}$ is mean Li–Yorke chaotic if and only if the following conditions hold:
  • There exist $m\in\mathbb{N}$ and an increasing sequence $(N_{k})_{k\in\mathbb{N}}$ of positive integers such that for each $k\in\mathbb{N}$ there are $r:=r(k)\in\mathbb{N}$ and scalars $b_{1,k},\cdots,b_{r,k}\in\mathbb{K}$ with $\left\|\sum_{j=1}^{r}b_{j,k}e_{j}\right\|_{p(k)}>0$ and
    $\frac{1}{N_{k}\left\|\sum_{j=1}^{r}b_{j,k}e_{j}\right\|_{p(k)}}\sum_{i=1}^{N_{k}}\left\|\sum_{j=1}^{r}b_{j,k}w_{j-i}\cdots w_{j-1}e_{j-i}\right\|_{m}\geq k,$
    where $p(k):=m$, if $1\leq k\leq m$, and $p(k):=k$, if $k>m$. Moreover, we consider $e_{k}:=(0)_{j\in\mathbb{N}}$ and $w_{k}=0$, for $k\leq 0$.
As a consequence of Theorem 31, we obtain the following corollary, which characterizes mean Li–Yorke chaos in the context of weighted backward shifts on Köthe sequence spaces, provided that these spaces satisfy condition (C). This corollary is obtained by directly applying Theorem 31 to the seminorms from Definition 6.
Corollary 34.
@@
    If $p=0$, then $B_{w}$ is mean Li–Yorke chaotic if and only if the condition (A) of Theorem 31 and the following condition hold:
    • (A1)
      There exist $m\in\mathbb{N}$ and an increasing sequence $(N_{k})_{k\in\mathbb{N}}$ of positive integers such that for each $k\in\mathbb{N}$ there are $r:=r(k)\in\mathbb{N}$ and scalars $b_{-r,k},\cdots,b_{r,k}\in\mathbb{K}$ with $\max_{-r\leq j\leq r}\left|a_{j,p(k)}b_{j,k}\right|\neq 0$ and
      $\frac{1}{N_{k}\max_{-r\leq j\leq r}\left|a_{j,p(k)}b_{j,k}\right|}\sum_{i=1}^{N_{k}}\max_{-r\leq j\leq r}\left|a_{j-i,m}b_{j,k}w_{j-i}\cdots w_{j-1}\right|\geq k,$
      where $p(k):=m$, if $1\leq k\leq m$, and $p(k):=k$, if $k>m$.

### 06 Apr 2026 v2

@@ 4 Mean Li–Yorke chaos
$\liminf_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d\big(f^{k}(x),f^{k}(y)\big)=0\quad\text{and}\quad\limsup_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d\big(f^{k}(x),f^{k}(y)\big)>0.$
The function $f$ is said to be mean Li–Yorke chaotic if there exists an uncountable set $S\subset X$ such that every pair $(x,y)\in S\times S$ of distinct points is a mean Li–Yorke pair.
We next present some definitions related to the notion of mean Li–Yorke chaos.
@@
Let $(X,d)$ be a metric space and let $f:X\to X$ be a continuous map.
@@
    We say that $f$ is mean sensitive if there exists $\delta>0$ such that, for every $x\in X$ and every $\varepsilon>0$, there exists $y\in X$ with $d(x,y)<\varepsilon$ and
    $\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(f^{i}(x),f^{i}(y))\geq\delta.$
  • (b)
    For a given positive number $\delta$, a pair $(x,y)\in X\times X$ is called a mean Li–Yorke $\delta$-chaotic pair if
    $\liminf_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(f^{i}(x),f^{i}(y))=0\quad\text{and}\quad\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(f^{i}(x),f^{i}(y))\geq\delta.$
    We say that $f$ is mean Li–Yorke sensitive if there exists $\delta>0$ such that, for every $x\in X$ and every $\varepsilon>0$, there exists $y\in X$ with $d(x,y)<\varepsilon$ such that $(x,y)$ is a mean Li–Yorke $\delta$-chaotic pair.
It is clear that every mean Li–Yorke sensitive system is mean sensitive. From now on, given a Fréchet space $X$ endowed with a family of seminorms $\bigl(\|\cdot\|_{n}\bigr)_{n\in\mathbb{N}}$, we will always consider the compatible metric $d$ to be the one given in (1). Recall that the compatible metric $d$ satisfies the following proprieties:
  • (P1)
    $d(x_{1}+x_{2},0)\leq d(x_{1},0)+d(x_{2},0)$ for all $x_{1},x_{2}\in X$.
  • (P2)
    $d(\lambda x_{1},0)\leq(1+|\lambda|)d(x_{1},0)$ for all $\lambda\in\mathbb{K}$ and $x_{1}\in X$.
To prove our results in this section, we need the following definitions and lemmas from [14].
Definition 23.
Let $X$ be a Fréchet space with the compatible metric $d$. Let $T:X\to X$ be a continuous linear operator. A vector $x\in X$ is called absolutely mean semi-irregular (or semi-irregular point) if
@@
Definition 24.
Let $T:X\to X$ be a continuous linear operator on a Fréchet space $X$ with the compatible metric $d$. The mean asymptotic cell and the mean proximal cell of 0 are defined by
$\text{MAsym}(T,0):=\left\{x\in X:\lim_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d(T^{k}x,0)=0\right\}\quad\text{and}$
$\text{MProx}(T,0):=\left\{x\in X:\liminf_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d(T^{k}x,0)=0\right\},\quad\text{respectively}.$
Remark 25.
As observed in [14], $\text{MProx}(T,0)$ is a $G_{\delta}$ set. Moreover, [14, Lemma 4.7] proves that if $\text{MAsym}(T,0)$ is residual, then $\text{MAsym}(T,0)$ coincides with the whole space.
Lemma 26.
[14, Proposition 4.11] Let $X$ be a Fréchet space, let $T:X\to X$ be a continuous linear operator, and let $d$ be a compatible metric on $X$. Then the following assertions are equivalent:
  • (a)
    $T$ is mean sensitive;
@@
    there exist a sequence $(y_{k})_{k}$ in $X$ and an increasing sequence $(N_{k})_{k}$ in $\mathbb{N}$ such that $\lim_{k\to\infty}y_{k}=0$ and
    $\inf_{k\in\mathbb{N}}\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}d(T^{i}y_{k},0)>0.$
@@
  • (a)
    $T$ is mean Li–Yorke chaotic;
  • (b)
    $T$ is mean Li–Yorke chaotic sensitive;
  • (c)
    $T$ admits an absolutely mean semi-irregular vector.
@@
[14, Theorem 4.27] Let $T:X\to X$ be a continuous linear operator on a Fréchet space $X$ with the compatible metric $d$. Then, the following statements are equivalent:
  1. (a)
@@
Let $X$ be a Fréchet sequence space over $\mathbb{Z}$, with the compatible metric $d$. Suppose that the sequence $(e_{n})_{n\in\mathbb{Z}}$ of canonical vectors is a basis. Suppose that the bilateral weighted backward shift $B_{w}$, with nonzero weights $w:=(w_{n})_{n\in\mathbb{Z}}$, is well-defined and continuous on $X$. Then the following statements are equivalent:
@@
    $\liminf_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d(w_{j-k}\cdots w_{j-1}e_{j-k},0)=0,$ for some $j\in\mathbb{Z}$;
@@
Proof.
It is obvious that (b) implies (a). Now, suppose that (a) holds for some $j\in\mathbb{Z}$. Then, there exists an increasing sequence $(n_{k})_{k\in\mathbb{N}}$ of positive integers such that
$\lim_{k\to\infty}\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}d(w_{j-i}\cdots w_{j-1}e_{j-i},0)=0.$
Take $\ell\in\mathbb{Z}$ with $\ell>j$ and define $N_{k}:=n_{k}+(\ell-j)$, for each $k\in\mathbb{N}$. Then
$\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}d(w_{\ell-i}\cdots w_{\ell-1}e_{\ell-i},0)$ $=\frac{1}{N_{k}}\sum_{i=1}^{\ell-j}d(w_{\ell-i}\cdots w_{\ell-1}e_{\ell-i},0)+\frac{1}{N_{k}}\sum_{i=\ell-j+1}^{N_{k}}d(w_{\ell-i}\cdots w_{\ell-1}e_{\ell-i},0)$
$\leq\frac{\ell-j}{N_{k}}+\frac{1}{N_{k}}\sum_{i=\ell-j+1}^{N_{k}}d(w_{\ell-i}\cdots w_{j-1}\overset{:=C}{\overbrace{w_{j}\cdots w_{\ell-1}}}e_{\ell-i},0)$
$=\frac{\ell-j}{N_{k}}+\frac{1}{N_{k}}\sum_{r=1}^{n_{k}}d(C\,w_{j-r}\cdots w_{j-1}e_{j-r},0)$
$\leq\frac{\ell-j}{N_{k}}+\frac{n_{k}(1+|C|)}{N_{k}}\frac{1}{n_{k}}\sum_{r=1}^{n_{k}}d(w_{j-r}\cdots w_{j-1}e_{j-r},0)\overset{k\to\infty}{\longrightarrow}0.$
Now, take $\ell\in\mathbb{Z}$ with $\ell<j$ and define $N_{k}:=n_{k}-(j-\ell)$, for each $k\in\mathbb{N}$. Suppose that $N_{k}>0$, for all $k\in\mathbb{N}$, otherwise, it suffices to discard the finitely many negative terms and reindex the sequence. Then
$\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}d(w_{\ell-i}\cdots w_{\ell-1}e_{\ell-i},0)$ $=\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}d(\overset{:=C}{\overbrace{\frac{1}{w_{\ell}\cdots w_{j-1}}}}w_{\ell-i}\cdots w_{\ell-1}w_{\ell}\cdots w_{j-1}e_{\ell-i},0)$
$\leq\frac{(1+|C|)}{N_{k}}\sum_{i=1}^{N_{k}}d(w_{\ell-i}\cdots w_{j-1}e_{\ell-i},0)$
$\leq\frac{(1+|C|)n_{k}}{N_{k}}\frac{1}{n_{k}}\sum_{r=(j-\ell)+1}^{n_{k}}d(w_{j-r}\cdots w_{j-1}e_{j-r},0)\overset{k\to\infty}{\longrightarrow}0.$
@@
Let $X$ be a Fréchet sequence space over $\mathbb{Z}$, endowed with the compatible metric $d$. Suppose that the sequence $(e_{n})_{n\in\mathbb{Z}}$ of canonical vectors is a basis and that condition (C) holds. Suppose that the bilateral weighted backward shift $B_{w}$, with nonzero weights $w:=(w_{n})_{n\in\mathbb{Z}}$, is well-defined and continuous on $X$. Then, $B_{w}$ is mean Li–Yorke chaotic if and only if the following conditions hold:
@@
    $\liminf_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d(w_{-k}\cdots w_{-1}e_{-k},0)=0;$
  • (B)
    there exist $\varepsilon>0$ and an increasing sequence $(N_{k})_{k\in\mathbb{N}}$ of positive integers, such that for each $k\in\mathbb{N}$ there are $r:=r(k)\in\mathbb{N}$ and scalars $b_{-r,k},\cdots,b_{r,k}\in\mathbb{K}$ with $d\left(\sum_{j=-r}^{r}b_{j,k}e_{j},0\right)<\frac{1}{k}$ and
    $\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}d\left(\sum_{j=-r}^{r}b_{j,k}w_{j-i}\cdots w_{j-1}e_{j-i},0\right)\geq\varepsilon.$
Proof.
($\Rightarrow$) Suppose that $B_{w}$ is mean Li–Yorke chaotic. Therefore, by Lemma 27, $B_{w}$ admits an absolutely mean semi-irregular vector $y:=(y_{j})_{j\in\mathbb{Z}}$, then
$\liminf_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d((B_{w})^{k}(y),0)=0\quad\text{and}\quad\limsup_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d((B_{w})^{k}(y),0)>0.$ (13)
Therefore, by Lemma 29, condition (A) holds. Since $B_{w}$ is mean Li–Yorke chaotic, then, by Lemma 27, $B_{w}$ is mean Li-Yorke sensitive and, in particular, it is mean sensitive. Thus, by Lemma 26, there exist $\varepsilon\in(0,1)$, a sequence $(z_{k})_{k}$ in $X$ and an increasing sequence $(N_{k})_{k}$ in $\mathbb{N}$ such that $\lim_{k\to\infty}z_{k}=0$ and
$\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}d((B_{w})^{i}(z_{k}),0)>\varepsilon,\quad\text{for all }k\in\mathbb{N}.$
Take an increasing sequence $(k_{s})_{s\in\mathbb{N}}$ of positive integers such that $d(z_{k_{s}},0)<\frac{1}{2s}$. Define $x_{s}:=z_{k_{s}}$ and $M_{s}:=N_{k_{s}}$ for each $s\in\mathbb{N}$. Then
$\frac{1}{M_{s}}\sum_{i=1}^{M_{s}}d((B_{w})^{i}(x_{s}),0)>\varepsilon,\quad\text{for all }s\in\mathbb{N}.$ (14)
Fix $s\in\mathbb{N}$ and write $x_{s}=(x_{s,n})_{n\in\mathbb{Z}}$. Since $\lim_{n\to\infty}\sum_{j=-n}^{n}x_{s,j}e_{j}=x_{s}$, then, by the continuity of $B_{w}$, there exists $r\in\mathbb{N}$ such that
$d\left((B_{w})^{i}\left(\sum_{j=-r}^{r}x_{s,j}e_{j}\right),(B_{w})^{i}(x_{s})\right)<\frac{\varepsilon}{2s},\quad\text{for all }i\in\left\{0,1,\cdots,M_{s}\right\}.$ (15)
Therefore, by (14) and (15), and using the triangle inequality of $d$ we obtain
$d\left(\sum_{j=-r}^{r}x_{s,j}e_{j},0\right)<\frac{1}{s}\quad\text{and}\quad\frac{1}{M_{s}}\sum_{i=1}^{M_{s}}d\left(\sum_{j=-r}^{r}b_{j,k}w_{j-i}\cdots w_{j-1}e_{j-i},0\right)>\frac{\varepsilon}{2}.$
($\Leftarrow$) By condition (A) and Lemma 29, we have that
If there exists $j_{0}\in\mathbb{Z}$ such that $\limsup_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d((B_{w})^{k}(e_{j_{0}}),0)>0$, then $B_{w}$ is mean Li–Yorke chaotic, given that $e_{j_{0}}$ is an absolutely mean semi-irregular vector. Otherwise, we have
Therefore, using properties (P1) and (P2) of the metric $d$, we get $c_{00}(\mathbb{Z})\subset\text{MProx}(B_{w},0)$. Therefore, $\text{MProx}(B_{w},0)$ is dense. By condition (B) and Lemma 26, we have that $B_{w}$ is mean sensitive. Then, there exists $x\in X$ such that
$\limsup_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d((B_{w})^{k}(x),0)>0.$
Thus, by Lemma 28, $B_{w}$ admits a dense set of absolutely mean semi-irregular vectors. Therefore, by Lemma 27, we conclude that $B_{w}$ is mean Li–Yorke chaotic. ∎
Remark 31.
$\lim_{n\to\infty}(B_{w})^{n}(e_{i})=0,\quad\text{for all }i\in\mathbb{N},$
Theorem 32.
Let $X$ be a Fréchet sequence space, endowed with the compatible metric $d$. Suppose that the sequence $(e_{n})_{n\in\mathbb{N}}$ of canonical vectors is a basis. Suppose that the unilateral weighted backward shift $B_{w}$, with nonzero weights $w:=(w_{n})_{n\in\mathbb{N}}$, is well-defined and continuous on $X$. Then $B_{w}$ is mean Li–Yorke chaotic if and only if the following condition holds:
  • There exist $\varepsilon>0$ and an increasing sequence $(N_{k})_{k\in\mathbb{N}}$ of positive integers, such that for each $k\in\mathbb{N}$ there are $r:=r(k)\in\mathbb{N}$ and scalars $b_{1,k},\cdots,b_{r,k}\in\mathbb{K}$ with $d\left(\sum_{j=1}^{r}b_{j,k}e_{j},0\right)<\frac{1}{k}$ and
    $\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}d\left(\sum_{j=1}^{r}b_{j,k}w_{j-i}\cdots w_{j-1}e_{j-i},0\right)\geq\varepsilon,$
    where we consider $e_{k}=(0)_{j\in\mathbb{N}}$ and $w_{k}=0$, for $k<1$.
As in the case of Li–Yorke chaos (see [10, Corollary 20]), we present below propositions that reveal a dichotomy of weighted shifts with respect to mean Li–Yorke chaos.
Proposition 33.
Let $X$ be a Fréchet sequence space over $\mathbb{Z}$, endowed with the compatible metric $d$. Suppose that the sequence $(e_{n})_{n\in\mathbb{Z}}$ of canonical vectors is a basis. Suppose that the bilateral weighted backward shift $B_{w}$, with nonzero weights $w:=(w_{n})_{n\in\mathbb{Z}}$, is well-defined and continuous on $X$. If $\liminf_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d(w_{-k}\cdots w_{-1}e_{-k},0)=0$, then either
  • (a)
    $B_{w}$ is mean Li–Yorke chaotic, or
  • (b)
    $\lim_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d((B_{w})^{k}(x),0)=0$, for all $x\in X$.
Proof.
Suppose that $B_{w}$ is not mean Li–Yorke chaotic. Then $\text{MProx}(T,0)=\text{MAsym}(T,0)$. Since
$\lim_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d((B_{w})^{k}(e_{0}),0)=\liminf_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d(w_{-k}\cdots w_{-1}e_{-k},0)=0,$
by Lemma 29 and proprieties (P1) and (P2) of the compatible metric $d$, we get $c_{00}(\mathbb{Z})\subset\text{MProx}(T,0)$. Therefore, by Remark 25, we have that $\text{MProx}(T,0)=\text{MAsym}(T,0)$ is residual. Thus, again by Remark 25, $\text{MAsym}(T,0)=X$. ∎
Proposition 34.
Let $X$ be a Fréchet sequence space, endowed with the compatible metric $d$. Suppose that the sequence $(e_{n})_{n\in\mathbb{N}}$ of canonical vectors is a basis. Suppose that the unilateral weighted backward shift $B_{w}$, with nonzero weights $w:=(w_{n})_{n\in\mathbb{N}}$, is well-defined and continuous on $X$. Then either
  • (a)
    $B_{w}$ admits a dense set of absolutely mean semi-irregular vectors, or
  • (b)
    $\lim_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d((B_{w})^{k}(x),0)=0$, for all $x\in X$.
Proof.
Suppose that item (b) is false. Then, there exists $x\in X$ such that
$\limsup_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d((B_{w})^{k}(x),0)>0.$
Since $\lim_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d((B_{w})^{k}(y),0)=0$, for all $y\in c_{00}(\mathbb{N})$, then, by Lemma 28, item (a) holds. ∎
By using Theorem 30, we obtain characterizations of mean Li-Yorke chaos in the context of weighted backward shifts on the Köthe sequence spaces $\lambda_{p}(A,\mathbb{Z})$ with $p\in[1,\infty)\cup\{0\}$, since these spaces satisfy condition (C). Furthermore, by using Theorem 32, we obtain the corresponding unilateral characterizations in this setting. Next, we will obtain results involving the increasing sequence of seminorms $(\|\cdot\|_{n})_{n\in\mathbb{N}}$ that induce the topology of $X$. To do this, we need the following definition and lemma:
Definition 35.
Let $X$ be a Fréchet space endowed with an increasing sequence $(\|\cdot\|_{k})_{k\in\mathbb{N}}$ of seminorms and with the compatible metric $d$. A vector $x\in X$ is called absolutely mean $m$-irregular if
Lemma 36.
Proposition 37.
Let $X$ be a Fréchet sequence space over $\mathbb{Z}$, endowed with an increasing sequence $(\left\|\cdot\right\|_{n})_{n\in\mathbb{N}}$ of seminorms and with the compatible metric $d$. Suppose that the sequence $(e_{n})_{n\in\mathbb{Z}}$ of canonical vectors is a basis and that condition (C) holds. Suppose that the bilateral weighted backward shift $B_{w}$, with nonzero weights $w:=(w_{n})_{n\in\mathbb{Z}}$, is well-defined and continuous on $X$. If $B_{w}$ is mean Li–Yorke chaotic, then the following conditions hold:
  • (A)
    $\liminf_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}d(w_{-k}\cdots w_{-1}e_{-k},0)=0;$
@@
The proof of item (A) is the same as the one given for item (A) in the proof of Theorem 30, so we only prove item (B). Since $B_{w}$ is mean Li-Yorke chaotic, then, by Lemma 27, $B_{w}$ admits an absolutely mean semi-irregular vector. Therefore, Lemma 36 ensures the existence of $m\in\mathbb{N}$ and $x=(x_{j})_{j\in\mathbb{Z}}\in X$ such that
@@
Thus, by (17) and (18), we conclude condition (B). ∎
We say that $X$ is a Banach sequence space over $\mathbb{Z}$ if $(X,\|\cdot\|)$ is a Banach space which is a vector subspace of $\mathbb{K}^{\mathbb{Z}}$ and the inclusion map $X\to\mathbb{K}^{\mathbb{Z}}$ is continuous, i.e., convergence in $X$ implies coordinatewise convergence. Under the assumption that $X$ is a Banach sequence space over $\mathbb{Z}$, we can improve Proposition 37. To do so, we will need the following lemma:
Lemma 38.
[7, Theorems 5 and 9] Let $(X,\|\cdot\|)$ be a Banach space and $T:X\to X$ a continuous linear operator. Then, the following assertions are equivalent:
  • (a)
    $T$ is mean Li–Yorke chaotic;
  • (b)
    $T$ admits an absolutely mean-irregular vector $x\in X$, that is,
    $\liminf_{N\to\infty}\frac{1}{N}\sum_{j=1}^{N}\|T^{j}x\|=0\quad\text{and}\quad\limsup_{N\to\infty}\frac{1}{N}\sum_{j=1}^{N}\|T^{j}x\|=\infty;$
  • (c)
    $T$ satisfies the Mean Li–Yorke Chaos Criterion (MLYCC), that is, there exists a subset $X_{0}$ of $X$ with the following properties:
    1. (i)
      $\liminf_{N\to\infty}\frac{1}{N}\sum_{j=1}^{N}\|T^{j}x\|=0$, for every $x\in X_{0}$;
    2. (ii)
      there are sequences $(y_{k})$ in $\overline{\operatorname{span}(X_{0})}$ and $(N_{k})$ in $\mathbb{N}$ such that
      $\frac{1}{N_{k}}\sum_{j=1}^{N_{k}}\|T^{j}y_{k}\|>k\|y_{k}\|,\quad\text{for every }k\in\mathbb{N}.$
@@
Let $(X,\|\cdot\|)$ be a Banach sequence space over $\mathbb{Z}$. Suppose that the sequence $(e_{n})_{n\in\mathbb{Z}}$ of canonical vectors is a basis and that condition (C) holds with the norm $\|\cdot\|$. Suppose that the bilateral weighted backward shift $B_{w}$, with nonzero weights $w:=(w_{n})_{n\in\mathbb{Z}}$, is well-defined and continuous on $X$. Then, $B_{w}$ is mean Li–Yorke chaotic if and only if the following conditions hold:
@@
    $\liminf_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}\|w_{-k}\cdots w_{-1}e_{-k}\|=0;$